OpenStudy (campayne):

In the following solution of 14t + 7 = 28, what is the reason for the third step? 1. 14t + 7 = 28 (Given) 2. 14t = 21 (Subtraction Property of Equality) 3. t = 3÷2 ( ? ) A) Substitution Property of Equality B) Division Property of Equality C) Multiplication Property of Equality D) Transitive Property of Equality

Mehek (mehek14):

\(\div=division\) so what answer choice goes with that?

OpenStudy (ciarán95):

In we're looking to try and solve for t, we need to get t on its own. So, given that we have 14t = 21 from the second step, we need to remove the '14' from in front of the t in some way. What so we do?? Well, what we can do is divide both sides by 14: \[\frac{ 14 }{ 14 }t = \frac{ 21 }{ 14 }\] 14/14 = 1, so we've solved our problem. Remember, we need to DIVIDE BOTH SIDES of the equation by the same thing, otherwise we are changing the equation itself, rather than just rearranging it. \[t = \frac{ 21 }{ 14 }\] We can simplify the fraction on the right by dividing top and bottom by 7 (the largest common factor): \[t = \frac{ 3 }{ 2 }\] So, given what we've just done, which of the options do you think we're proving @CamPayne ?

OpenStudy (campayne):

well it can't seem to be anything else but option B

OpenStudy (ciarán95):

Yes, that's correct! :)